# Green Operations of SWIPT Networks: The Role of End-User Devices

### v1

@@ III-B Base Station and UE Energy Consumption Model
The power consumed by a BS depends on several factors, which vary according to the BS type (e.g. macro, micro, femto) and the implementation technology (e.g. standalone vs cloud-RAN), among others. In what follows, we adopt a very flexible BS energy model, first proposed in [31, 32]:
$q_{1}+U_{d}[q_{2}+q_{3}(P-P_{min})]$ (3)
where $U_{d}$ is the downlink BS utilization and $P$ is its transmit power, which we assume varies within the interval $[P_{min},P_{max}]$. $q_{1}$ models the total power consumed when the BS is idle, which does not depend on load or transmit power. In what follows we assume that, when transmitting, a BS is using the whole bandwidth at its disposal. $q_{2}U_{d}$ models that fraction of consumed power which depends on utilization, but not on transmit power (such as part of baseband and RF processing). The third component models the fraction of consumed power due to the power amplifier which depends, at the same time, on the transmit power $P$ and on the amount of active resources at the BS.
@@ III-C Base station service model
We define the utilization $U(S(x))$ of the BS serving a UE at location $x$ as the average fraction of time in which the BS is busy transmitting (in downlink) or receiving (in uplink), respectively. Thus, the expression of the utilization of the BS is given by the fraction of BS time dedicated to all active users:
@@
IV Modeling User Perceived Performance
In this section, we characterize the main performance parameters, e.g. the per-bit delays and the harvested power perceived by a typical user who is just beginning service [^2] This assumption is standard in works exploiting stochastic geometry. In fact there is no difference with respect of the case where users reside in the system since a while, it is just a way of choosing the point of view from which to carry out the calculations. , as well as the mean harvested power, as a function of the main system parameters. The expression of the power harvested by a user at $x$ is given by the following result.
@@ IV Modeling User Perceived Performance
$I(x)$ is the total power harvested by the user at $x$ from BSs other than the one with which it is associated. $O(x)$ is the power harvested from UE transmissions, averaged over time, and $U_{d}(S(x))$ is the downlink utilization of the BS with which the user in $x$ is associated.
@@
For IoT users, the ideal per-bit delay perceived in uplink is $\tau_{u,I}^{id}(x)=\delta_{u}\tau_{u}^{id}(x)$. With $\bar{\tau}_{j}$ ($\bar{\tau}_{j,I}$), $j=d,u$ we denote the average per-bit delay perceived in downlink (resp. in uplink) by broadband and IoT users, respectively. The following result derives an expression for the average per-bit delay perceived by the typical user (BB or IoT) which is just beginning service.
@@
$A(r,x,\theta)$ is the area of the circle centered at $(x,\theta)$ with radius $x$ that is not overlapped by the circle centered at $(0,-r)$ with radius $r$. $C(r,P,G,\bar{I})$ is given by (2), with the interference term $\bar{I}$ given by
@@
For the proof, please refer to appendix -B. We say that the considered system is in the dense IoT regime when the density of IoT users is such that the probability of having a cell without IoT users is negligible. This assumption is coherent with many current projections on expected IoT deployments in 6G, in which the expected device densities are of the order of 10 million devices/km2 and more [34, 35, 36, 37].
@@
(Maximally fair GPS weights). The values of the GPS weight $w_{d}$ which maximize fairness among users are given by $w_{d}=\delta_{d}(1-\eta)$ for the TS mode, and
@@
for the PS modes. $\bar{r}$ is the mean distance of UEs from the serving base station, and $\bar{I}(\bar{r},k)$ is given by Theorem 1.
@@
However such a ratio is a function of $r$, and of the interference experienced by the specific user considered. Thus, we consider the mean user, and set $w_{d}$ in such a way as to satisfy the equality
@@ V Energy-Optimal Network Configuration
One of the main open issues in SWIPT networks is to determine, as a function of the main system parameters as well as of the energy and traffic demands, how the main network parameters should be tuned in order to optimize the energy consumed by the system. To this end, we present below the formulation of the optimization problem which provides, for a given BS energy model, as well as for a given user mean density, the energy optimal BS transmit power, the optimal density of active BSs, as well as the optimal amount of BS time dedicated to power transfer, which satisfy the performance constraints.
@@
(TS version)
$min_{P,\lambda_{b},\eta}\;\;\lambda_{b}\left[q_{1}+\frac{\bar{\tau}_{d}(\eta,P,\lambda_{b})}{\tau^{0}_{d}}\left(q_{2}+q_{3}(P-P_{min})\right)\right]$
@@
$\frac{\bar{\tau}_{d}(\eta,P,\lambda_{b})}{\tau^{0}_{d}}\leq 1,\quad$ $\frac{\bar{\tau}_{u}(\eta,P,\lambda_{b})}{\tau^{0}_{u}}\leq 1$ (20)
$P_{min}\leq P\leq$ $P_{max}$ (21)
$0\leq\eta\leq$ 11 11 (22)
$CDF_{h}(h_{0},\eta,P)\leq$ $\mu$ (23)
$0\leq\lambda_{b}\leq$ $\lambda_{b,max}$ (24)
where $h_{0}$ in constraint (23) is the minimum harvested power required by each IoT device to operate, computed based on the amount of energy required by the device during a whole on-off cycle. $\mu$ is the maximum acceptable ratio of IoT users which harvest less than $h_{0}$ Joules per second. $\bar{\tau}_{d}$, $\bar{\tau}_{u}$ are given by Theorem 1. $\lambda_{b,max}$ derives from practical constraints to BS deployments in urban settings. In the PS configurations, Problem 1 takes a very similar structure, where the optimization variable $\nu$ takes the place of variable $\eta$, with $0\leq\nu\leq 1$. In both formulations, Problem 1, is clearly non-convex and nonlinear, and it cannot be solved efficiently.
A simple heuristic approach to solve Problem 1 is based on defining, for each of the three decision variables, a discrete set of values (e.g., based on practical considerations on the maximal accuracy with which they can be varied in practical scenarios), and on an exhaustive grid search on those sets. However, such a brute-force approach gets quickly unfeasible when the decision variables are varied with fine granularity. Thus, for such large problems, a metaheuristic is required [38]. To this end, we have adopted a Genetic Algorithm (GA) approach [39], where a population of chromosomes, each constituting the binary representation of a triplet of values for the three decision variables, is evaluated by assigning a fitness value. Such a value is a score indicating the optimality of the solution, and it is used to feed the archive of the best solutions found so far, by picking up the dominant one (i.e., the chromosome with the highest fitness value, with respect to our problem). Then a new population is generated using proper genetic operators (Gaussian Mutation and Crossover, which for each couple of chromosomes returns a child that lies halfway on the segment defined by the two parents) for the mating and mutations of the randomly selected chromosomes. Such a sequence of steps is repeated until a termination condition (i.e. on the maximum number of iterations) is met. In this way, guided by the fitness function, the GA realizes a random exploration of the solution space, based on the evolutionary operators’ output.
@@
To facilitate the convergence of the GA, the fitness function has been defined through the penalty method [40], by which, in addition to the cost function of 1, a penalty is introduced, proportional to the degree by which constraints (20) (downlink and uplink) and (22) are violated. This ensures that the metaheuristic is able to explore values of the decision variables that are very close to the border of the feasibility region of Problem 1. The resulting fitness function has the following expression:
$v(P,\lambda_{b},\eta)=\lambda_{b}\left(q_{1}+\frac{\bar{\tau}_{d}(\eta,P,\lambda_{b})}{\tau^{0}_{d}}\right)[q_{2}+q_{3}(P-P_{min})]+k_{1}\left(\frac{\bar{\tau}_{u}(\eta,P,\lambda_{b})}{\tau^{0}_{u}}-1\right)^{+}-k_{2}[CDF_{h}(h_{0},\eta,P)-\mu]^{-}+$
$+k_{3}\left(\frac{\bar{\tau}_{d}(\eta,P,\lambda_{b})}{\tau^{0}_{d}}-1\right)^{+}$
where $(x)^{+}=max(x,0)$, and $(x)^{-}=min(x,0)$. The (positive) values of $k_{1}$, $k_{2}$, and $k_{3}$ may be tuned empirically to ensure the desired behavior of the GA metaheuristic in proximity to the border of the feasibility region.
@@ VI-A Setup
We assumed base stations work at a frequency of $1.5$ GHz, and use a bandwidth of $50$ MHz, compatible with 4G+ standards. Unless otherwise indicated, we assume a percentage of IoT devices equal to $80\%$ of the total number of UEs (representative of many present-day scenarios), a transmit power equal to $0.2$ W for both IoT and BB UEs, and a frequency reuse factor of $3$. Moreover, by default, we assumed a beamforming gain equal to $10$, and constant over the whole main lobe aperture (which we assume to be equal to $45$ degrees), and a path loss exponent equal to $3$, typical of urban areas. For linear energy harvesting, we assumed a conversion efficiency of $0.9$, with no lower/upper threshold. We assume the deployed base stations to be of the macro type, with a transmit power that varies between $1$ and $11$ W. Unless otherwise stated, we set a target mean per-bit delay in downlink for BB (resp. IoT) UEs equal to $10^{-5}$ s (resp. $10^{-3}$ s), and in uplink equal to $10^{-4}$ s for all UEs, (e.g., typical of IoT systems for environmental monitoring [15]). We consider the user density to vary from $10^{-4}$ users per $m^{2}$ (modeling settings with a high share of BB users) to $10^{-1}$ users per $m^{2}$ (modeling scenarios with crowds of BB UEs and high-density IoT deployments). Unless otherwise specified, we assume IoT UEs to be active all the time ($\phi=1$). We set to $5\%$ the maximum acceptable share of IoT users who are not able to harvest the target minimum energy.
@@
The parameters of the BS energy model are chosen to fit two different types of BSs. The first type (labeled LLP – low load proportionality) reflects the behavior of the majority of current stand-alone BSs, and it is characterized by a $27\%$ load proportionality (with $q_{1}=1100$, $q_{2}=100$, and $q_{3}=30$). Conversely, the high load proportionality (HLP) BS type (with $q_{1}=482.3$, $q_{2}=48.23$, and $q_{3}=144.69$) corresponds to a $75\%$ load proportionality, achievable, e.g., through time-domain duty-cycling at the sub-system level, i.e., through micro-sleep techniques involving modules of the BS or of the BBU in cloud-RAN designs [41]. For the sake of comparison, these parameters were chosen to fit a per-BS maximum consumed power of $1500$ W, typical of stand-alone macro BSs [42]. The value of BS density has been varied with a granularity of $10^{-4}$ $m^{-2}$, to guarantee good accuracy of the GA search process. We assumed the mean number of users per base station to be lower bound to $5$. This models the simple energy-saving strategy common among MNOs, which switch off those BSs that serve very few users to no user at all, as they represent a very high energy cost per user and a small benefit for performance. The size of the initial population in the GA algorithm has been set to $100$ chromosomes. This choice has proven to be a good compromise between computational load and convergence speed. In the fitness function, parameters $k_{1}$, $k_{2}$ and $k_{3}$ have been set to $1,000$, $5,000$, and $1,000$, respectively, as these values have proven appropriate to enable our GA approach to explore effectively the borders of the feasibility region. The termination condition of the GA has been set based on the convergence of the fitness value. Specifically, we assumed that convergence is reached when the geometric average of the relative change in the value of the spread over $50$ generations is less than $10^{-6}$, and the final spread is less than the mean spread over the past $50$ generations.
VI-B Linear Energy Harvesting

### v3

@@ III-B Base Station and UE Energy Consumption Model
The power consumed by a BS depends on several factors, which vary according to the BS type (e.g. macro, micro, femto) and the implementation technology (e.g. standalone vs cloud-RAN), among others. In what follows, we adopt a very flexible BS energy model, first proposed in [31, 32], by which the power consumed by a BS, denoted as $P_{BS}$, is given by the following expression:
$P_{BS}=q_{1}+U_{d}[q_{2}+q_{3}(P-P_{min})]$ (3)
In this model, the power consumed is thus given by the sum of three contributions.
  • A first contribution, given by $q_{1}$, is constant, and depends only on the specific type of BS considered (e.g. macro, micro, femto, cloud). It models the power consumed when the BS is not carrying any traffic to/from users. As such, it does not depend on utilization or transmit power, and it is due to the power consumed by, e.g., part of the cooling function, by power amplifier consumption in idle state, and by all those functions which keep the BS in an operational (i.e. non standby) state.
  • The second component $q_{2}U_{d}$, in which $U_{d}$ is the downlink BS utilization, models that fraction of consumed power which depends on BS utilization (and it is thus proportional to the amount of traffic served), but not on transmit power. It models the power consumed by such functions as baseband processing and RF signal processing.
  • Finally, the third component $q_{3}U_{d}(P-P_{min})$ (where and $P$ is its transmit power, which we assume varies within the interval $[P_{min},P_{max}]$) models the fraction of consumed power due to the power amplifier which depends, at the same time, on the transmit power $P$ and on the fraction of time that the BS is busy transmitting (given by $U_{d}$).
This energy model is very flexible and suited for accounting not only for implementations with different degrees of load proportionality, but also for cloud RAN configurations.
@@ III-C Base station service model
We define the utilization $U(S(x))$ of the BS serving a UE at location $x$ as the average fraction of time in which the BS is busy transmitting (in downlink – $U(S_{d}(x))$) or receiving (in uplink – $U(S_{u}(x))$), respectively. Thus, the expression of the utilization of the BS is given by the fraction of BS time dedicated to all active users:
@@
IV Modeling User-Level Performance
In this section, we characterize the main performance parameters, i.e., the per-bit delay and the harvested power of a typical user who is just beginning service [^2] This assumption is standard in stochastic geometry, and it is one of the possible ways of choosing the point of view from which to carry out the analysis. Indeed, there is no difference in terms of performance with respect of the case in which users have been in the system for a while. , as well as the mean harvested power, as a function of the main system parameters. The expression of the power harvested by a user at $x$ is given by the following result.
@@ IV Modeling User-Level Performance
$I(x)$ is the total power harvested by the user at $x$ from BSs other than the one with which it is associated. $O(x)$ is the power harvested from UE transmissions, averaged over time, and $U_{d}(S(x))$ is the downlink utilization of the BS with which the user at $x$ is associated.
@@
For IoT users, the ideal per-bit delay perceived in uplink is $\tau_{u,I}^{id}(x)=\delta_{u}\tau_{u}^{id}(x)$. With $\bar{\tau}_{j}$ ($\bar{\tau}_{j,I}$), $j=d,u$ we denote the average per-bit delay perceived in downlink (resp. in uplink) by broadband and IoT users, respectively. By leveraging stochastic geometry, we now derive the main analytical results, in terms of probabilistic expressions for user-perceived performance metrics, which account for the random spatial patterns of both BSs and UEs. The following result derives an expression for the average per-bit delay perceived by the typical user (BB or IoT) which is just beginning service.
@@
$A(r,x,\theta)$ is given by $A(r,x,\theta)=\pi x^{2}-\bigg{[}r^{2}\arccos\left(\frac{r+x\sin(\theta)}{d(r,x,\theta)}\right)+x^{2}\arccos\left(\frac{x+r\sin(\theta)}{d(r,x,\theta)}\right)+\bigg{.}$
.$\bigg{.}-\frac{1}{2}\sqrt{[r^{2}-(d(r,x,\theta)-x)^{2}][(d(r,x,\theta)+x)^{2}-r^{2}]}\bigg{]}$, and $d(r,x,\theta)$ is the euclidean distance between $(x,\theta)$ and $(0,-r)$. $C(r,P,G,\bar{I})$ is given by (2), with the interference term $\bar{I}$ given by
@@
For the proof, please refer to appendix -B. We say that the considered system is in the dense IoT regime when the density of IoT users is such that the probability of having a cell without IoT users is negligible. This assumption is coherent with many current projections on expected IoT deployments in 6G, in which the expected device densities are of the order of 10 million devices per km2 and more [34, 35, 36, 37].
@@
(Maximally fair GPS weights). The values of the GPS weight $w_{d}$ which maximize fairness among users are given by $w_{d}=\delta_{d}(1-\eta)$ for the time split (TS) mode, and
@@
for the power split modes. $\bar{r}$ is the mean distance of UEs from the serving base station, and $\bar{I}(\bar{r},k)$ is given by Theorem 1.
@@
However such a ratio is a function of $r$, and of the interference experienced by the specific considered user. Thus, we consider the mean user, and set $w_{d}$ in such a way as to satisfy the equality
@@ V Energy-Optimal Network Configuration
One of the main open issues in SWIPT networks is to determine, as a function of the main system characteristics as well as of the energy and traffic demands, how the main network parameters should be tuned in order to optimize the energy consumed by the system. To this end, in this section we elaborate the formulation of the optimization problem which provides, for a given BS energy model, as well as for a given user mean density, the energy optimal BS transmit power, the optimal density of active BSs, as well as the optimal amount of BS time dedicated to power transfer, which satisfy the specified performance constraints.
The objective function is given by the mean power per unit surface consumed by the network, as a function of the BS density, of the BS transmit power, and of the time/power switch ratio. It is thus derived as the product of the power consumed by a BS (expressed in Equation (3)) and of the mean BS density. Note that, by the definition of per bit delay, $U_{d}=\frac{\bar{\tau}_{d}(\eta,P,\lambda_{b})}{\tau^{0}_{d}}$. The optimization problem which can be formulated with such an objective function clearly depends also on the EH operating mode. In the TS case, we have the following optimization problem formulation:
@@
(TS operating mode)
$\underset{P,\lambda_{b},\eta}{\text{minimize}}\ \lambda_{b}\left[q_{1}+\frac{\bar{\tau}_{d}(\eta,P,\lambda_{b})}{\tau^{0}_{d}}\left(q_{2}+q_{3}(P-P_{min})\right)\right]$ (20)
@@
  1. (C1)
    $\frac{\bar{\tau}_{d}(\eta,P,\lambda_{b})}{\tau^{0}_{d}}\leq 1,\quad\frac{\bar{\tau}_{u}(\eta,P,\lambda_{b})}{\tau^{0}_{u}}\leq 1$
  2. (C2)
    $P_{\text{min}}\leq P\leq P_{\text{max}}$
  3. (C3)
    $0\leq\eta\leq 1$
  4. (C4)
    $CDF_{h}(h_{0},\eta,P)\leq\mu$
  5. (C5)
    $0\leq\lambda_{b}\leq\lambda_{b,\text{max}}$
Constraints C1 require the mean BS utilization in downlink and uplink to be smaller than one. The expression of the ideal per-bit delay is the one from Theorem 1. Constraints C2 and C3 define the range of acceptable values for the BS transmit power and the time split ratio, respectively. Constraint C4 requires that the probability for an IoT device to harvest less power than the minimum required to operate (denoted as $h_{0}$, and accounting for the amount of energy required by the device during a whole on-off cycle) be less than a given maximum acceptable value $\mu$. The expression of the cumulative distribution function is the one derived in Theorem 2. Finally, constraint C5 derives from practical limitations to the maximum density of BS deployments in realistic urban settings, which in any realistic scenario cannot exceed a maximum value.
In the power splitting EH operating modes (both static and dynamic), the resulting optimization problem takes a very similar structure. Constraint C3 is replaced by
[…]
The other constraints and the objective function remain the same, except that the formulas for $CDF_{h}$ and $\tau_{d}$ are those specific to the SPS and DPS modes.
In all formulations, however, the optimization problem is non-convex and nonlinear, and it cannot be solved efficiently. This is since $\bar{\tau}_{d}$ (which appears in the expression of both the objective function and in constraints C1) is a nonlinear and non-convex function of $P$, $\lambda_{b}$, and $\eta$ (or $\mu$). Similar consideration holds also for the expression of $CDF_{h}(h_{0},\eta,P)$ from Theorem 2, which is nonlinear and non-convex in the three optimization variables.
V-A GA Metaheuristic
One of the possible heuristic approaches to solving Problem 1 is based on an exhaustive grid search and on defining a discrete set of values (e.g., based on practical considerations regarding the maximal accuracy with which they can be varied in practical scenarios) for each of the three decision variables. However, such a brute-force approach becomes quickly unfeasible when the decision variables must be quantized with fine granularity. In this cases, a metaheuristic is required [38]. Among these, those based on Genetic Algorithms (GA) [39] are particularly well suited for their ability to avoid getting trapped in local minima by maintaining a population of potential solutions. In addition, they are amenable to parallel implementations, and despite not being tied to a specific problem structure, they can be adapted to the specific problem for better efficiency. Indeed, as represented in Algorithm 1, the GA considers a population of $n$ chromosomes, each constituting the binary representation of a triplet of values for the three decision variables, is evaluated by assigning the value associated with a fitness function. Such a value is a score indicating the optimality of the solution, and it is used to feed the archive of the best solutions found at a given point in time by the heuristic. This is implemented by picking up the dominant one (i.e., the chromosome with the highest fitness value). Then, a new population is generated using proper genetic operators. For the selection, we leveraged classical stochastic uniform selection [40], which preserves diversity in the population and helps avoiding premature convergence to suboptimal solutions. Then, we utilize Gaussian Mutation [41] as the default mutation operator, while the Crossover Heuristic [40] mechanism generates offsprings positioned along the line connecting the two parents at a distance of $ratio$, favoring the direction away from the parent with lower fitness.
Algorithm 1 Genetic Algorithm for Solving Problem 1
1:Initialize the population $W$ with $n$ chromosomes, where each chromosome represents a binary-encoded triplet $(P,\lambda_{b},\eta)$ corresponding to the decision variables;
2:repeat
3:     Evaluate the fitness of each chromosome in $W$ using the penalty-based fitness function (Eq. 22);
4:     Select parents for mating based on their fitness values using stochastic uniform selection;
5:     Generate an offspring by using Crossover Heuristic;
6:     Mutate the offspring using Gaussian Mutation;
7:     Evaluate the new offspring and integrate them into the population $W$ by replacing the least fit individuals;
8:until Iterations = = $maxGen$ OR no improvement in the best fitness value for $i$ consecutive iterations.
The algorithm iterates for a maximum of $maxGen$ generations or stops earlier if the termination condition is met, defined as no improvement in the best fitness value for $i$ consecutive generations. In this way, guided by the fitness function, the GA realizes an optimistic exploration of the solution space based on the evolutionary operators’ output.
@@ V-A GA Metaheuristic
To facilitate the convergence of the GA, the fitness function is defined through the penalty method [42], by which, in addition to the cost function of Problem 1, a penalty is introduced as a linear combination of the constraints, in such a way that the fitness function value increases proportionally to the degree by which constraints (C1) and (C3) over the utilization and minimum harvested power are violated. This ensures that the resulting metaheuristic can also explore those values of the decision variables that are very close to the border of the feasibility region of Problem 1. The resulting expression for the fitness function is as follows:
$v(P,\lambda_{b},\eta)=$ $\lambda_{b}\left\{q_{1}+\frac{\bar{\tau}_{d}(\eta,P,\lambda_{b})}{\tau^{0}_{d}}[q_{2}+q_{3}(P-P_{min})]\right\}+k_{1}\left(\frac{\bar{\tau}_{u}(\eta,P,\lambda_{b})}{\tau^{0}_{u}}-1\right)^{+}+$ (22)
-- $k_{2}[CDF_{h}(h_{0},\eta,P)-\mu]^{-}+k_{3}\left(\frac{\bar{\tau}_{d}(\eta,P,\lambda_{b})}{\tau^{0}_{d}}-1\right)^{+}$
where $(x)^{+}=max(x,0)$, and $(x)^{-}=min(x,0)$. $k_{1}$, $k_{2}$, and $k_{3}$ are nonnegative weights, assigned empirically to facilitate the efficient exploration of the border of the feasibility region. In our experiments, we have verified that by choosing for the three weights a value of the same order of magnitude as the maximum value of the objective function in Problem 1, i.e. equal to $\lambda_{b,max}[q_{1}+q_{2}+q_{3}(P_{max}-P_{min})]$ is sufficient to significantly enhance both convergence consistency and solution quality with respect to the approach without penalty, effectively accelerating the convergence of the algorithm.
@@ VI-A Setup
We assume base stations work in the $1.5$ GHz band and use a bandwidth of $50$ MHz, compatible with 4G+ standards. Unless otherwise indicated, we assume a percentage of IoT devices equal to $80\%$ of the total number of UEs (representative of many present-day scenarios), a transmit power equal to $0.2$ W for both IoT and BB UEs, and a frequency reuse factor of $3$. Moreover, by default, we assume a beamforming gain equal to $10$, constant over the whole main lobe aperture (which we assume to be equal to $45$ degrees), and a path loss exponent equal to $3$, typical of urban areas. For linear energy harvesting, we assume a conversion efficiency of $0.9$, with no lower/upper threshold. We assume the deployed base stations to be of the macro type, with a transmit power that varies between $1$ and $11$ W. Unless otherwise stated, we set a target mean per-bit delay in downlink for BB (resp. IoT) UEs equal to $10^{-5}$ s (resp. $10^{-3}$ s), and in uplink equal to $10^{-4}$ s for all UEs, (e.g., typical of IoT systems for environmental monitoring [15]). We consider the user density to vary from $10^{-4}$ users per $m^{2}$ (modeling settings with a high share of BB users) to $10^{-1}$ users per $m^{2}$ (modeling scenarios with crowds of BB UEs and high-density IoT deployments). Unless otherwise specified, we assume IoT UEs to be active all the time ($\phi=1$). We set to $5\%$ the maximum acceptable share of IoT users who are not able to harvest the target minimum energy.
@@
The parameters of the BS energy model are chosen to fit two different types of BSs. The first type (labeled LLP – low load proportionality) reflects the behavior of the majority of current stand-alone BSs, and is characterized by a $27\%$ load proportionality (with $q_{1}=1100$, $q_{2}=100$, and $q_{3}=30$). Conversely, the high load proportionality (HLP) BS type (with $q_{1}=482.3$, $q_{2}=48.23$, and $q_{3}=144.69$) corresponds to a $75\%$ load proportionality, achievable, e.g., through time-domain duty-cycling at the sub-system level, i.e., through micro-sleep techniques involving modules of the BS or of the BBU in cloud-RAN designs [43]. For the sake of comparison, these parameters were chosen to fit a per-BS maximum consumed power of $1500$ W, typical of stand-alone macro BSs [44]. The value of BS density has been varied with a granularity of $10^{-4}$ $m^{-2}$, to guarantee good accuracy of the GA search process. We assume the mean number of users per base station to be lower bounded by $5$. This models the simple energy-saving strategy common among MNOs, which switch off those BSs that serve very few users to no user at all, as they represent a very high energy cost per user and a small benefit for performance. The size $n$ of the initial population in the GA algorithm was set to $100$ chromosomes. This choice proved to be a good compromise between computational load and convergence speed. In the fitness function, parameters $k_{1}$, $k_{2}$ and $k_{3}$ were set to $1,000$, $5,000$, and $1,000$, respectively, as these values proved appropriate to enable our GA approach to explore effectively the borders of the feasibility region. The termination condition of the GA was set based on the convergence of the fitness value. Specifically, we assume that convergence is reached when the geometric average of the relative change in the value of the spread over ($i=10$) generations is less than $10^{-6}$, and the final spread is less than the mean spread over the past $i$ generations. We also set a generation limit $m$ to 150 to prevent potential infinite iterations. This value was chosen based on an empirical evaluation of the typical number of iterations required across different starting points. Regarding the Crossover Heuristic, we choose a $ratio=1.9$, to have a good distance from parents, and prevent the algorithm from easily getting stuck at a local minimum.
VI-B Computational Complexity
As for the time complexity of the proposed GA, in the worst case scenario, if $n$ iterations are required per each population of $m$ chromosomes, and if with $T_{sim}$ we denote the time to compute the value of the function modeled by the chromosome and to perform the estimation of the fitness value, the time complexity of our algorithm is $\mathcal{O}(mnT_{sim}$). In our setup, consisting in an Intel Core i5 6-core - used in parallel - processor and 16 GB 2667 MHz DDR4 RAM, $T_{sim}$ has been always less than 2s. However, we have empirically observed that in $x$ tries of the GA, out of the number of points evaluated for all the six configurations analyzed, our termination condition always brought the GA to converge before the maximum number of iterations. This confirmed the validity of our heuristic choice of the three constants $k_{1}$, $k_{2}$, and $k_{3}$, as it is hard to model analytically the mutual relationships between these constants and their impact on the resulting value assumed by the overall function.
VI-C Linear Energy Harvesting